# Remote Rowhammer Attack using Adversarial Observations on Federated Learning Clients

### v1

@@ IV-D Bit-Flip Validation
IV-D1 Estimated Row Activation Counts
Table V summarizes the calculated expected row activations for all evaluated ASR models at the medium sparsity setting (0.1%), comparing our two-stage PPO method (PPO-TS) against single-stage PPO (PPO-SS) and Unperturbed baseline methods.
TABLE V: Estimated Row Activations Based on Empirical Clustering (Top 0.05% Sparsity Level)
Model Method $H_{\text{max}}(0.0005)$ RUR (%) $\mathbb{E}_{Activations}$
Conformer-CTC-S PPO-TS 68.3 202,168
PPO-SS 296,000 53.9 159,544
Unperturbed 9.3 27,528
QuartzNet 5x5 PPO-TS 71.1 273,024
PPO-SS 384,000 57.2 219,648
Unperturbed 8.7 33,408
Citrinet 256 PPO-TS 66.4 166,000
PPO-SS 250,000 52.9 132,250
Unperturbed 10.2 25,500
Squeezeformer-XS PPO-TS 68.1 194,766
PPO-SS 286,000 54.4 155,584
Unperturbed 8.1 23,166
Eff. Conformer-CTC PPO-TS 66.1 129,556
PPO-SS 196,000 56.9 111,524
Unperturbed 9.6 18,816
IV-D2 Estimated Bit Flips
Refer to Table II, averaging bit flip thresholds over different single sided aggressor patterns, bit flips are expected when repeated activation exceeds 240K, with more vulnerable cells flips at 200K, i.e., $\overline{T_{\text{flip}}}=240,000$. Our agent’s attack on QuartzNet 5x5 with sparsity of 0.05% can be deemed successful in inducing Rowhammer bit flips, and there is a high possibility of success in inducing Rowhammer bit flips through attacking Conformer-CTC-S.
V Discussion
@@
V-A Practical Implications and Feasibility
Our experiments provide empirical evidence that PPO-generated physical-domain perturbations can effectively induce clustered and repetitive parameter updates critical for successful Rowhammer exploitation in federated learning scenarios. Particularly at higher sparsity levels in large-scale FL deployments, our PPO-TS approach significantly outperforms baseline methods by achieving high repetition rates of over 70% and highly localized cluster densities. Such performance substantially increases the likelihood of repeatedly activating specific DRAM rows, thus enhancing the practical feasibility of successful Rowhammer bit-flips.
@@ V-A Practical Implications and Feasibility
V-B Limitations and Open Challenges
We acknowledge two primary limitations affecting the practical applicability of the proposed PPO-driven Rowhammer attack:
@@ V-B Limitations and Open Challenges
    Lack of Control over Specific Victim Rows: The proposed indirect Rowhammer attack inherently lacks precise control over the specific DRAM rows affected by induced bit-flips, making outcomes unpredictable and potentially transient. This uncertainty reduces the reliability of achieving specific attack objectives (e.g., targeted data corruption or privilege escalation). Future research should explore advanced side-channel techniques or hardware-software co-design approaches to improve the precision and predictability of the Rowhammer effects induced by physical-domain perturbations.
V-C Security Considerations for Federated Learning Systems
Our findings highlight significant security risks arising from commonly used performance optimization techniques in federated learning systems, including sparse updates, RDMA, huge pages, and pinned memory. While these optimizations greatly enhance computational and communication efficiency, they also inadvertently create predictable and sustained DRAM activation patterns exploitable by hardware-level adversaries. Consequently, FL practitioners must carefully reconsider optimization-related trade-offs, explicitly evaluating hardware security vulnerabilities when designing and deploying federated learning systems, especially in sensitive or critical infrastructure scenarios.
V-D Potential Mitigation Strategies
Based on our comprehensive analysis and findings, we propose several practical defensive measures to mitigate the identified threat:
@@ V-D Potential Mitigation Strategies
  • Systematic Analysis of Optimization Technique Deployment: When deploying new optimization techniques, systematically assess and analyze the effects of new techniques in combination of existing optimization techniques with considerations of both software execution and hardware characteristics, hence avoiding inherent security risks and lowering the feasibility of the proposed attack vector.
V-E Broader Impact and Future Directions
Our study demonstrates a compelling need for interdisciplinary approaches combining hardware security, adversarial machine learning, and federated learning systems design. By integrating hardware-level vulnerabilities into FL threat modeling, we advocate a holistic view that proactively identifies, assesses, and mitigates emerging threats.
Future research directions include:
@@ V-E Broader Impact and Future Directions
VI Conclusion
In this paper, we introduced and comprehensively evaluated a novel Rowhammer attack vector targeting federated learning (FL) systems, exploiting physically driven interference orchestrated by a Proximal Policy Optimization (PPO)-based reinforcement learning agent. Unlike traditional Rowhammer attacks, our method leverages adversarial acoustic or electromagnetic perturbations at client sensors, indirectly inducing clustered and repetitive parameter updates on FL servers. Our empirical results demonstrate the efficacy of our approach, particularly at higher sparsity levels common in realistic FL deployments, effectively increasing the frequency of repeated DRAM row activations critical for triggering Rowhammer bit-flips.
Through extensive evaluation and analysis across diverse Automatic Speech Recognition (ASR) architectures, we demonstrated that the proposed two-stage PPO approach substantially enhances adversarial capabilities compared to baseline methods, achieving both higher repeated update rates and more concentrated update clustering. Our joint evaluation highlighted practical feasibility, with achieved activation counts realistically approaching Rowhammer thresholds under favorable conditions, emphasizing the tangible security threat posed by this attack vector.
@@ VI Conclusion
Our findings underscore the necessity for interdisciplinary threat modeling, encompassing both software and hardware domains, to ensure the secure and dependable deployment of federated learning systems.

### v2

@@ V-B Bit-Flip Validation
V-B1 Estimated Row Activation Counts
Table III summarizes the calculated expected row activations for evaluated models at both sparsity settings. Although the QuartzNet 5x5 attack achieve slight higher RUR, the larger size and higher precision limits the $H_{\text{max}}$, and therefore, lower $\mathbb{E}_{Act}$. Among these target models, Conformer-CTC-S and MobileNetV3 have the higher $\mathbb{E}_{Act}$, both above 200K under the two sparsity settings.
V-B2 Estimated Bit Flips
Refer to Table II, averaging bit-flip thresholds over different single-sided aggressor patterns, bit flips are expected when repeated activation exceeds 240K, i.e., $\overline{T_{\text{flip}}}=240,000$. Under the 0.1% sparsity setting, MobileNetV3 is the only model whose estimated activation count exceeds the average threshold, indicating that Rowhammer bit flips can be reliably induced. Conformer-CTC-S remains close to the vulnerable range and may still induce bit flips under favourable conditions, whereas Squeezeformer-XS and QuartzNet 5x5 fall below the practical threshold and are therefore unlikely to succeed. Under the 0.05% sparsity setting, the attack becomes more feasible overall. Conformer-CTC-S, Squeezeformer-XS, and MobileNetV3 all exceed the average threshold, while QuartzNet 5x5 reaches 233K activations, placing it between the minimum and average threshold and indicating that bit flips can still be induced, although with greater difficulty than other three models.
VI Discussion
@@
VI-A Practical Implications and Feasibility
Our experiments provide empirical evidence that PPO-generated physical-domain perturbations can effectively induce clustered and repetitive parameter updates critical for successful Rowhammer exploitation in federated learning scenarios.
@@ VI-A Practical Implications and Feasibility
VI-B Limitations and Open Challenges
We acknowledge two primary limitations on the practical applicability of the proposed remote Rowhammer attack:
@@ VI-B Limitations and Open Challenges
    Lack of Control over Specific Victim Rows: The proposed indirect Rowhammer attack inherently lacks precise control over the specific DRAM rows affected by induced bit-flips, making outcomes unpredictable and potentially transient. This uncertainty reduces the reliability of achieving specific objectives (e.g., targeted data corruption or privilege escalation). Future research should explore advanced side-channel techniques or hardware-software co-design approaches to improve the precision and predictability of the Rowhammer effects induced by physical-domain perturbations.
VI-C Mitigation Measures
Based on our findings, we propose several practical defensive measures to mitigate the identified threat:
@@ VI-C Mitigation Measures
  • Systematic Security Analysis on Deployment: on deployment of FL systems, all optimization techniques, software or hardware based ones, should be systematically assessed across both software execution and hardware behavior, to prevent unintentional susceptibility to threat vectors such as our proposed vector when improving computational and communication efficiency. To reduce the feasibility of the proposed attack vector, practitioners should evaluate optimization trade-offs not only from a performance perspective but also with respect to their security implications, particularly in sensitive or critical infrastructure environments.
VI-D Broader Impact and Future Directions
Our study demonstrates a need for interdisciplinary approaches combining hardware security, adversarial machine learning, and federated learning systems design. By integrating hardware-level vulnerabilities into FL threat modelling, we advocate a holistic view that proactively identifies, assesses, and mitigates emerging threats, leading to the following future research directions:
@@ VI-D Broader Impact and Future Directions
VII Conclusion
In this paper, we introduced and comprehensively evaluated a novel Rowhammer attack vector targeting federated learning (FL) systems, exploiting physically driven interference orchestrated by a Proximal Policy Optimization (PPO)-based reinforcement learning agent. Unlike traditional Rowhammer attacks, our method leverages adversarial acoustic or electromagnetic perturbations at client sensors, indirectly inducing clustered and repetitive parameter updates on FL servers. Our empirical results demonstrate the feasibility of our propose attack vector, effectively increasing the frequency of repeated DRAM row activations for triggering Rowhammer bit-flips.
Through extensive evaluation and analysis across diverse Automatic Speech Recognition (ASR) architectures as well as a CIFAR10 classification MobileNetV3, we demonstrated that the proposed PPO approach consistently achieves repeated clustered updates. The evaluation highlighted practical feasibility, with achieved activation counts realistically approaching Rowhammer thresholds under favorable conditions, emphasizing the tangible security threat posed by this attack vector.
@@ VII Conclusion
Our findings underscore the necessity for interdisciplinary threat modelling, encompassing both software and hardware domains, to ensure the secure and dependable deployment of federated learning systems.